Find the minimum initial speed of a cork that pops out of a bottle and which travels a horizontal distance of 10 m.
1
Expert's answer
2020-04-09T09:25:13-0400

The range for the projectile motion

"R=\frac{v_0^2\sin2\theta}{g}."

The maximum range for the projectile motion ("\theta=\pi/4" )

"R_{\rm max}=\frac{v_0^2}{g}."

Hence

"v_0=\sqrt{gR_{\rm max}}=\sqrt{10\times 10}=10\:\rm m/s."